# Multiwavelength Observations of the Apparently Non-repeating FRB 20250316A

### v1

@@ 5 Summary
In this paper, we present multiwavelength follow-up observations of the nearby bright apparent one-off FRB 20250316A, including radio observations with FAST, X-ray observations with EP and Chandra, and optical observations with WFST and SVOM/VT. The non-detection of pulses with the FAST radio telescope suggests that this burst is different from the current repeating FRB sample, likely a one-off event. The EP-WXT obtained the first upper limit of at just 3.5 hours post-burst, followed by the EP-FXT achieving a much deeper limit of at 1.4 days after the FRB. The subsequent Chandra observations, 11 days post-burst, established the most constraining upper limit on the flux of , corresponding to an X-ray luminosity upper limit of in the 0.5–10 keV band, lower than most ULXs. These results place one of the most stringent limits on persistent X-ray emission from a non-repeating FRB, disfavoring the association of FRBs with ULXs and providing critical constraints for theoretical models of FRB progenitors and emission processes. In addition, the combination of the optical and X-ray upper limits constrain the kinetic energy of afterglow models to .

### v2

@@ 5 Summary
In this paper, we present multiwavelength follow-up observations of the nearby bright apparent one-off FRB 20250316A, including radio observations with FAST, X-ray observations with EP and Chandra, and optical observations with WFST and SVOM/VT. The non-detection of pulses with the FAST radio telescope indicates that the predicted rate of well-studied repeating FRBs at the FAST limiting fluence is – orders of magnitude higher than FAST’s 3 upper limit, and that the effective FRB rate at the CHIME limiting fluence is lower than the rates of all known repeating FRBs. These results suggest that this burst is distinct from the current sample of repeating FRBs and is likely a one-off event, supporting the results of CHIME/FRB Collaboration et al. 2025. The EP-WXT obtained the first upper limit of at just 3.5 hours post-burst, followed by the EP-FXT achieving a much deeper limit of at 1.4 days after the FRB. The subsequent Chandra observations established the most constraining 0.5–10 keV flux upper limit of , corresponding to an X-ray luminosity upper limit of , which is lower than those of most ULXs. These results place one of the most stringent limits on persistent X-ray emission from a non-repeating FRB, disfavoring the association of FRBs with ULXs and providing critical constraints for theoretical models of FRB progenitors and emission processes. In addition, the combination of the optical and X-ray upper limits constrain the kinetic energy of afterglow models to .